OpenStudy (anonymous):

I need help understanding how to write an equation for this worded problem... the relationship between the amount of time a zebra runs at maximum speed and the distance it covers is shown... time (min) 3,6,9,12,15 distance (mi) 2,4,6,8,10 write an equation to describe this relationship I need help understanding how to write an equation for this worded problem... the relationship between the amount of time a zebra runs at maximum speed and the distance it covers is shown... time (min) 3,6,9,12,15 distance (mi) 2,4,6,8,10 write an equation to describe this relationship @Mathematics

OpenStudy (saifoo.khan):

\[x+\frac x2\]

OpenStudy (anonymous):

That doesn't work?

OpenStudy (anonymous):

x = 3,6,9,12,15 y=2,4,6,8,10

OpenStudy (saifoo.khan):

\[x = y + \frac y2\]

OpenStudy (saifoo.khan):

Check, \[x = 10 + \frac{10}{2}\]\[x = 10+5\]x = 15

OpenStudy (radar):

Or d=t/1.5

OpenStudy (radar):

Distance (d) equals time (t) divided by 1.5

OpenStudy (anonymous):

Thanks! Both of you!

OpenStudy (radar):

U R welcome, good luck with your studies.